OpenStudy (anonymous):

Find the polynomial function with roots 1, 7, and -3 of multiplicity 2. (1 point) Help?? Anyone?!!!?!?!

OpenStudy (anonymous):

\[(x-1)(x-7)(x+3)^2\] or whatever you get when you multiply that out

OpenStudy (anonymous):

What would be the three of them multiplied? I've only ever multiplied two.

OpenStudy (anonymous):

actually 4

OpenStudy (anonymous):

\[(x-1)(x-7)(x+3)^2=(x-1)(x-7)(x+3)(x+3)\] etc

OpenStudy (anonymous):

How do I multiply 4 together? Do I distribute x from the first one to every single one and so on?

OpenStudy (amistre64):

just work them 2 at a time: distribution is a good notion to keep in mind yes \[(x-1)(x-7)(x+3)(x+3)\] \[a(x-7)(x+3)(x+3)\] \[(ax-7a)(x+3)(x+3)\] \[b(x+3)(x+3)\] \[(bx+3b)(x+3)\] \[c(x+3)\] \[cx+3c\]

OpenStudy (amistre64):

working backwards ... \[cx+3c~:~\text{but c=bx+3b}\] \[(bx+3b)x+3(bx+3b)\] \[bx^2+3bx+3bx+9b\] \[bx^2+6bx+9b~:~\text{but b=ax-7a}\] \[x^2(ax-7a)+6x(ax-7a)+9(ax-7a)\] \[ax^3-7ax^2+6ax^2-42ax+9ax-64a\] \[ax^3-ax^2-33ax-64a\] ...but a = x-1